OpenStudy (anonymous):

Help with 6th grade math

OpenStudy (anonymous):

What is our question?

OpenStudy (anonymous):

ok one second please

OpenStudy (anonymous):

take your time

OpenStudy (anonymous):

regular price = $100; discount rate = 35%

OpenStudy (anonymous):

And you're looking for the price after the discount right?

OpenStudy (anonymous):

Find the discount

OpenStudy (anonymous):

I have some more after this;'(

OpenStudy (anonymous):

well if you have a calculator handy this is an extremely simple question... the equation should be... 100x.35=?

OpenStudy (anonymous):

3500

OpenStudy (anonymous):

so 35

OpenStudy (anonymous):

yup! That's how much is being taken away from the original price, so you're saving $35

OpenStudy (anonymous):

another

OpenStudy (anonymous):

go ahead, i've got about another 15 minutes

OpenStudy (anonymous):

The regular price of an item is $120. The item is on sale at a discount rate of 20%. What is the sale price of the item?

OpenStudy (anonymous):

It's the same thing. you always take the original number and multiply by the percentage. the way to get your percentage in your equation correctly is to make it a decimal. meaning in this case you'd make 20% .20

OpenStudy (anonymous):

I got 2400

OpenStudy (anonymous):

nope.. you're forgetting the decimal point. it's 120 X .20 = ??? if you don't make the %20 a decimal number you'll end up way higher then you need to be and get the question wrong

OpenStudy (anonymous):

24

OpenStudy (anonymous):

yup that's the discount... now if you want to find the price of the item after you take that discount away you would subtract that discount from your original price. in this case it would be 120-24=???

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

another

OpenStudy (anonymous):

The regular price of an item is $65. The item is on sale with a discount of $8. What is the discount rate

OpenStudy (anonymous):

That is the problem

OpenStudy (anonymous):

sorry let me catch up here. for your previous question whatever you get from the 120-24 is your final answer yes... and sadly i'm out of time so you'll need to have the next person come along and help.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

goodluck!

OpenStudy (anonymous):

the last question I asked is wht I need the answer to

OpenStudy (anonymous):

The regular price of an item is $65. The item is on sale with a discount of $8. What is the discount rate?

OpenStudy (mathstudent55):

Ok, $65 item. Discount of $8 What percent was the discount?

OpenStudy (mathstudent55):

You want to know $8 is what percent of $65. To find what percent a number is of another number, divide the part by the total and multiply by 100.

OpenStudy (mathstudent55):

$8 is what percent of $65: $8/$65 * 100

OpenStudy (anonymous):

812.5

OpenStudy (anonymous):

I do not get the answer:'(

OpenStudy (anonymous):

and I have one more question

OpenStudy (mathstudent55):

8 divided by 65 = 0.123 0.123 * 100 = 12.3 Answer 12.3%

OpenStudy (mathstudent55):

Ok, go for it.

OpenStudy (anonymous):

What is the discount rate for a $120 camera that is on sale for $105?

OpenStudy (mathstudent55):

First find the discount. Not the percent discount. Just the discount. What is 120 - 105?

OpenStudy (anonymous):

15

OpenStudy (mathstudent55):

Great. Now we want to know what percent 15 is of 120.

OpenStudy (mathstudent55):

Divide the part by the whole and multiply by 100. 15/120 * 100 = 0.125 * 100 = 12.5%

OpenStudy (anonymous):

one more sorry

OpenStudy (mathstudent55):

Ok, last one bec gtg.

OpenStudy (mathstudent55):

Are you still there?